OpenStudy (anonymous):

log base 16 of 49 - log base 4 of 7. Explain

hero (hero):

Use the change of base formula to simplify this

OpenStudy (anonymous):

\[(\log49/\log16)-(\log7/\log4)\] Then what?

hero (hero):

You have to use change of base formula appropriately to simplify this. Hint: \[\log_{7} 7^2 = 2\]

OpenStudy (anonymous):

But there are no bases that work with 49 and 16 or 7 and 4. You're left with: \[\log_{4}49/2 - \log_{2}7/4 \]

hero (hero):

\[\frac{\log_4{49}}{\log_4{4^2}} = \frac{1}{2}\log_4{49}\] Now you have: \(\dfrac{1}{2}\log_4{7^2} - \log_4{7} = \dfrac{2}{2}\log_4{7} - \log_4{7} \) \( = 0\)

hero (hero):

Do you get it now?

OpenStudy (anonymous):

Sort of... The general sense. Thanks.

OpenStudy (anonymous):

All set!